OpenStudy (anonymous):

Find, correct to 2 decimal places, the angle between the tangents to y =3e^-x and y =2+ e^x at their point of intersection.

OpenStudy (amistre64):

well, where do they meet at? that point might be good to know; and then we can derive these at with the point to find the tangents vector

OpenStudy (amistre64):

2+e^x = 3e^(-x) wolfram might help to find the intersection

OpenStudy (amistre64):

(0,3), that was nice

OpenStudy (amistre64):

soo, [3e^(-x)]' = -3e^(-x) at x=0; slope = -3 [2+e^x]' = e^x at x=0; slope = 1

OpenStudy (amistre64):

the vectors are then <1,-3> and <1,1>, we can dot these to find the angle

OpenStudy (amistre64):

1.1 + -3.1 -2 -2 -------------- = ------- = ------- sqrt(2) * sqrt(10) sqrt(20) 2sqrt(5) cos(a) = -1/sqrt(5); a = cos-1(-1/sqrt(5)) if I did it right

OpenStudy (amistre64):

116.57 maybe?

OpenStudy (amistre64):

or; one angle is 45 the other is: tan(a) = -3 a = tan-1(-3) = |-71.565....| a+45 = 116.565.....

OpenStudy (amistre64):

same get up

OpenStudy (anonymous):

the answer is 63.43 I dont know how they got it

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=y%3D2%2Be^x+and+y%3D3e^%28-x%29 this the graph so I know the point is good; I know the derives are good, so it must be in how im interpreting the angles

OpenStudy (anonymous):

i got that far also

OpenStudy (amistre64):

slope of 1 = 45 degrees; maybe they want the radians?

OpenStudy (amistre64):

no matter how i do it I only see an angle of 116.57 degrees; you sure your answer is to this problem?

OpenStudy (amistre64):

180 - 116.57 = 6.43

OpenStudy (amistre64):

they wanted the supplement angle to the one we found

OpenStudy (amistre64):

seeing how there are 2 angles formed by lines that cross; one of ours was 116.57 and the other was 63.43

OpenStudy (anonymous):

ok cool

OpenStudy (amistre64):

how they expect you to decide is beyond me lol

OpenStudy (anonymous):

its assumed to be the acute angle

OpenStudy (anonymous):

the way I learnt this way by memorising a formula\[\tan(\theta) = |\frac{m1-m2}{1+(m1)(m2)}|\]

OpenStudy (anonymous):

where m1 and m2 are the gradients of the curves at the point of interest , I did use vectors or anything like that. The absolute value in the formula gives a acute angle for the answer